Can you show me the steps for this equation.:
5y-4x=29 6y-5x=35

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

---
5y-4x=29
6y-5x=35
---
put the system of linear equations into standard form:
---
-4x+5y=29
-5x+6y=35
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = -1
y = 5
---
now by elimination:
---
( -4x+5y=29 ) * -5
( -5x+6y=35 ) * 4
---
20x-25y=-145
-20x+24y=140
---
add the system:
---
-y=-5
y = 5
---
-4x+5(5)=29
-4x+25=29
-4x=4
x = -1
---
solution by elimination method:
x = -1
y = 5
---
both method agree !
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by Alan3354(69443)

Can you show me the steps for this equation.:
5y-4x=29 6y-5x=35
------------
There are 2 equations.
It's a system of 2 equations.
----
Step 1, say what you want to do.